# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v4

@@ 6.2.4. Knowledge Base Updation
Out of the 91 responses edited by experts, 72 (79.1%) received approval from the knowledge base expert (D4) to be added to the knowledge base. For 49.3% of these, the bot’s final response (generated by merging the initial response and expert edit suggestion) underwent additional editing by the knowledge base expert. These changes were minimal with an average relative Hamming distance of 28.2%, and mainly addressed aspects of “tone and structure”, such as preferring active voice and making responses more generally applicable across patients. This hints that the bot effectively incorporated experts’ edits.
Experts noticed improvements in the completeness and accuracy of CataractBot’s generated answers over time, thus reducing their workload: “With time, I think the answers did get better… I am mostly just saying yes.” (D3). However, there was an instance where the knowledge base expert approved incorrect information (i.e., the hospital’s phone number with an extra digit) to be added to the knowledge base, which CataractBot provided to three information seekers, before it was fixed. Despite expanding the knowledge base, we identified our bot’s knowledge to be limited in two medical areas—lens types and their comparisons, and lab tests—and one logistical area—insurance tie-ups. Addressing these required manually adding new documents.
6.3. Privacy and Trust
@@ 6.3.1. Personalization vs Privacy
On the other hand, fueled by the recent hype around AI and LLMs, eight participants expected “precise” and “personalized” answers to their queries with little concern for privacy, and were slightly disappointed with generic responses. This aligns with prior findings indicating minimal privacy concerns related to digital data in the global south, specifically in India (Okolo et al., 2021). Interestingly, three participants did not even consider their medical data to be “private”, as they stated, “It’s okay, it’s nothing personal to be hidden. All (of us) have common health issues.” (A14). Moreover, two participants questioned why others would be interested in their medical data, as “it is not useful for anyone else” (P18). Two participants wanted personalization, along with the sharing of patient data, to be an “opt-in” feature where individuals conduct a cost-benefit analysis and make an informed choice based on their preferences. They believed that a personalized bot would offer more relevant, and fewer IDK, responses.
@@
“If you don’t want to share your information [with the bot], it should say “If you ask personal questions, I won’t have access, so you may have to visit hospital.”… We have to compromise.” – A21.
@@
Five experts voted to integrate the chatbot with patient data, including schedules and consultation history, to offer personalized responses. D3 suggested collecting basic patient medical history to facilitate customized responses from the bot and doctors: “We don’t even know whether they’re diabetic, they’re hypertensive. So, when they ask “What should I eat and come?”… I ain’t sure what to tell.” (D3). Although the onboarding form attempts to capture such information in the ‘Extra Details’ field, none of our participants provided that information due to time constraints. However, D1 and D2 also voiced skepticism towards the LLMs capability to understand complex patient records accurately. “There are a lot of variables… the surgeon, the patient, their tests… It’s too much information to process.” (D1). In addition, the knowledge base expert cautioned that highly personalized responses would result in minimal updates to the knowledge base, potentially increasing CataractBot-related workload for experts and negatively impacting chatbot adoption among experts.
CataractBot operates as a user-driven platform from the perspective of patients and attendants, while it functions as a bot-driven system for experts. Information seekers initiate conversations by asking questions and receive responses. In contrast, the bot initiates and drives verification-related conversations with experts. Patients and attendants appreciated that CataractBot “does not disturb” them, and provides information solely based on the questions asked. However, a few participants suggested the bot should proactively offer relevant information.
6.3.2. Reliability and Trust
Six information seekers explicitly mentioned trusting the bot as a system because of their trust in the hospital, where they had chosen to undergo surgery. As CataractBot was deployed through Sankara Eye Hospital’s verified Facebook Business account, its default WhatsApp username included the hospital’s name and displayed a green tick indicating ‘verified’ status (Figures 1, 2). Moreover, the integration of the bot with the hospital was reinforced by the patient coordinator advising them to register and use the bot.
Participants gained trust in the bot’s responses due to the human experts verifying answers (Figure 1F). As A12 stated: “Basically the trust only came when I saw the green tick mark… Before that, I was also in the question mark zone… So I will say it’s a machine… But then somebody is checking it, confirming it… It’s actually then trustworthy.”. The experts-in-the-loop ensured that the bot’s response were more reliable and “not fake” (P18). This is in contrast to a few participants who complained about receiving false information from Google search results.
Although the experts-in-the-loop feature increased trust in final responses, it had a negative impact on perceived bot intelligence. Four participants noted that when an expert corrected CataractBot’s initial answer, trust in the bot’s unverified answers reduced. For instance, “After some time, when the doctor said it is invalid, I was like, okay, should I even trust the bot?” (P19). In total, 29.4% of the bot’s generated answers for medical and logistical queries were marked incorrect by the expert. This could be attributed to the limited custom knowledge base, the experts’ high standards for the tone and structure of answers, or our prompting strategy favoring caution. A9 pointed: “It [The bot] said that cough is dangerous on the surgery day as it increases eye pressure. But the doctor said it is completely fine… [The bot] unnecessarily alarmed me.”
6.3.3. Accountability
CataractBot provides a persistent written record of expert-patient communication, which has implications. Patients mentioned utilizing the information provided by CataractBot as a reference and sharing it with others undergoing cataract surgery. Also, this information can be valuable when discussing their surgery with experts outside of the Sankara Eye Hospital ecosystem: “[Patients] can share the same information with their doctor at their native place… So that [communication with another doctor] becomes a little easier.” (D1). However, it also increased the sense of accountability among experts, making them “somewhat nervous”. They felt the need to be “very lawyer-like in our conversation because we don’t want it to come back and bite us tomorrow.” (D4). Furthermore, coordinators identified the system’s potential to identify logistical bottlenecks in the hospital. For example, if multiple patients are inquiring about their surgery time, it may indicate a lapse in timely communication, leading them to reach out to the bot.
@@ 7. Discussion
In this paper, we propose CataractBot, an LLM-powered experts-in-the-loop WhatsApp chatbot to address the information needs of patients undergoing cataract surgery. Our user study, investigating how different features of CataractBot affected its usage and integration into patient-provider workflows, revealed positive evidence of its usability among the different stakeholders. Doctor-verified responses were key to prompting patients’ and attendants’ engagement with the bot, and establishing their trust. Patients and attendants appreciated that CataractBot instantly answered questions anytime, reducing the need for hospital calls or visits and saving time. Those with limited (tech-)literacy found it easy to use because of the multilingual and multimodal support. Doctors and patient coordinators commended CataractBot for acting as a facilitator, creating a layer of privacy between them and their patients, and providing them the flexibility to verify responses at their convenience.
We designed and evaluated our experts-in-the-loop chatbot system in the context of cataract surgery, as the straightforward surgical procedure, brief duration and high success rate provided an ideal low-stakes pilot environment. Moving forward, our open-sourced framework can be extended to assist with complex, atypical, or long-term treatments–e.g., glaucoma surgery, cancer treatment, or pregnancy–where easy access to verified information would be valuable for patients. The applicability can be extended further, to fields beyond the medical domain, such as law, finance and education. In these scenarios, end-users (e.g., defendants, taxpayers, or students) could receive synchronous responses from an LLM using a custom knowledge base, and subsequent asynchronous verified responses from experts (e.g., lawyers, accountants, or teachers). Expert-mediated bots can even be used to upskill lesser-trained members of the workforce as well. For instance, these bots could serve as a valuable learning resource for community health workers, who typically have minimal (medical) training. The resource could boost their confidence, while helping them provide accurate information to their care recipients.
@@ 7.1.1. Expert-in-the-loop Approaches
Prior studies have raised concerns about the application of LLMs in healthcare scenarios, due to their inconsistency, potential errors, and bias (Denecke et al., 2024; Au Yeung et al., 2023; Bérubé et al., 2021). Moreover, users who are less (technologically) literate face a heightened risk of harm (Kapania et al., 2022). Our experts-in-the-loop framework not only addresses these concerns but also adheres to OpenAI’s usage policies (OpenAI, 2024), which state that “tailored medical/health advice cannot be provided without review by a qualified professional”. These chatbots operate at the “sweet spot of patient-LLM-clinician collaboration” (Hao et al., 2024), where the capabilities of technology and human judgment are integrated. In our study, this ensured reliability and cultivated trust, as information seekers knew that their operating doctor was verifying responses. However, trust in the initial automated answers reduced when these answers were subsequently marked as incorrect/incomplete by the doctor. To address this in future work, one design approach could be to eliminate the initial automated response, providing only expert-verified answers. We caution that this might result in long wait times, which could discourage bot usage and prompt users to seek information elsewhere. Nonetheless, for high-risk use cases, such as cancer treatment, this approach may be appropriate, as users may value verified accuracy over immediacy. Future research could involve an empirical comparison of two paradigms: a hybrid system with synchronous automated answers followed by asynchronous verification versus a fully asynchronous system delivering only verified answers. Such a study would provide valuable insights into balancing immediate access with the accuracy required for critical medical information in chatbot design.
7.1.2. Reduce Workload for Experts
Deploying CataractBot in the hospital increased the workload of doctors due to verification and knowledge base updates. Future implementations should simplify experts’ tasks. To minimize their verification effort, one approach is to deliver pre-verified answers to specific questions that the expert has repeatedly marked as correct. Given the potential for LLM hallucinations, cached responses must be carefully matched to similar patient demographics and question phrasing before reuse. Apart from that, an information management tool would help support efficient knowledge base updates. For CataractBot, we collaborated with hospital staff to curate a custom knowledge base; however, during deployment, we observed experts having to frequently add information to the bot’s responses, highlighting the limitations of our initial corpus. For longitudinal deployments of such bots, we recommend a user-friendly system for experts to review and expand the knowledge base without technical assistance—such as adding new documents to a designated folder. As the knowledge base evolves, a mechanism is needed to periodically assess the quality of LLM-generated answers. Gumma et al. (2024) found that GPT-4’s evaluations of AI-generated responses align well with human expert assessments. Incorporating an evaluator LLM within the framework would automatically ensure that the bot’s response quality remains intact, so that experts’ workload indeed reduces as its knowledge base expands.
7.1.3. Improve Support for Multimodal Communication
@@
Our system faced challenges in accurately interpreting questions from voice messages due to limitations in transcription and translation technologies. The same inaccurate transcriptions were shared with doctors leading to confusion. Additionally, non-verbal cues in audio messages–such as signs of anxiety, impatience, or the presence of others–were lost in textual transcriptions. This hindered experts’ ability to deliver appropriate care (also noted in (Li et al., 2023)). To address this, we suggest sharing the original audio clips with experts. Recognizing that having to listen to each audio message would be added workload for experts, the textual transcription should be shared upfront, with the option to access the audio on-demand. Beyond audio, considering the visual capabilities of recent LLM models like GPT-4V, such bots could be enhanced to process image-based inputs, aligning with recommendations in prior studies (Jain et al., 2018). For example, a patient could upload a photo of a specific medicine to ask questions about it, instead of providing a written description. While this could improve the user experience for information seekers, it may increase the workload for experts. Further, prior research (Hao et al., 2024) cautions that although LLMs are proficient in text processing, their ability to handle other forms of media, such as image and video, remains less reliable.
7.1.4. Personalization
@@
While personalized responses can boost adoption and utility by providing more relevant information, they also raise privacy concerns, potentially reducing usage among certain users. Additionally, experts may face increased workloads as they reference user-specific information for verification, since many personalized edits may not be suitable for inclusion in the generic knowledge base. Low (tech-)literate users might require assistance in making informed decisions about the use of their data (Okolo et al., 2021). A binary approach of either sharing all records or none may not address individual preferences. Instead, users should have the agency to selectively share records they deem relevant and acceptable for the bot’s access. This aligns with the philosophy of India’s National Digital Health Mission (National Health Authority and Family Welfare, 2020). One solution could be to allow users to upload relevant documents or manually enter specific details they want the bot to access. This can empower users to control the information shared with the bot while benefiting from a more personalized experience.
@@ 7.2. Limitations
We acknowledge several limitations of this work. First, results should be interpreted with caution, given the specificities of the study, including a relatively small sample size, users’ first encounter with a medical chatbot, and the cataract treatment scenario. Positive user responses, minimal increases in expert workload, and other findings should be validated in future studies, preferably with longitudinal designs. Second, system-level analyses, such as scalability testing and power measurement, are needed before deploying the proposed solution more broadly. Finally, although participants reported high trust, repeated IDK responses or inaccurate information during regular use may erode this trust. We note the potential risk of disseminating inaccurate medical information at scale with chatbot and LLM technologies. Therefore, stronger methods are needed to develop, review, and update the knowledge base, as well as to routinely evaluate the accuracy of bot’s responses.

### v5

@@ 6.2.6. Knowledge Base
Out of the 91 responses edited by experts, 72 (79.1%) received approval from the knowledge base expert (D4) to be added to the knowledge base. For 49.3% of these, the bot’s final response (generated by merging the initial response and expert edit suggestion) underwent additional editing by the knowledge base expert. These changes were minimal with an average relative Hamming distance of 28.2%, and mainly addressed aspects of “tone and structure”, such as preferring active voice and making responses more generally applicable across patients. This hints that the bot effectively incorporated experts’ edits. Between the first and sixth weeks of the study, the proportion of LLM-generated answers that experts marked as “accurate and complete” increased by 9.9% for medical questions and 22.9% for logistical questions. This suggests that as the knowledge base was updated, the experts’ verification workload decreased.
6.3. Bot Integration into Doctor-Patient Workflows
6.3.1. Workflows
The majority of messages (71.4%) were sent in the 8-day period from 5 days prior to surgery to 2 days post-surgery with the highest number of messages (76) sent on the surgery day (Figure 3B). Messages were sent throughout the day (7am to 11pm), as A18 mentioned: “Whenever I am idle at work or at home… whatever (question) comes to my mind, I just ask.” The workflow employed by information seekers varied. Most participants would ask a question, leave their phones, and check the answer later when free. They hardly noticed the wait due to being occupied with their day-to-day activities. In such cases, notifications played a key role. The message stating that the expert had provided a verified response (Figure 1G), made the otherwise unnoticeable green tick (Figure 1F) prominent. Participants, when in need of a quick response, either relied on the initial (LLM-generated) answer or re-verified it on the internet.
19/22 survey respondents (86.4%) agreed that the bot responded quickly. On average, the bot took 11.810.2 seconds to respond. Expert verification took 162.9172.3 minutes. We found verification without edits (151.1147.1 minutes) to be significantly faster than verification with edits (191.3172.4 minutes) with t(265)=2.7, p¡0.05. Participants were “okay to wait” for the expert’s verification, since the response was being verified by “busy doctors”. This demonstrated their clear understanding of the bot’s workflow.
Notifications played a key role. The message stating that the expert had provided a verified response (Figure 1G), made the otherwise unnoticeable green tick (Figure 1F) prominent and the bot’s process self-explanatory.
For experts, the asynchronous nature of the bot provided them the flexibility to offer verifications at their convenience. The majority of responses (31) were provided by experts at 4-5 pm, post their busy workday, as most OPD and surgeries finish by 4 pm. The second-highest responses (28) occurred at 1-2 pm, during their lunch break. This usage pattern was described as: “I use my mobile phone only when I go for breaks… lunch break, coffee break, or in the morning before I start OPD. The rest of the time, my (cellular) data is off.” (D1). Experts refrained from using their phones in front of patients due to the negative perception associated with using phones for personal reasons. We found that the experts were “fine” with the additional task of verifying the bot’s responses, as articulated by D4: “It takes only about 15-20 seconds to answer a question… It was not really eating into my time… It’s much easier than replying to my mailbox.”
As five experts used CataractBot on their personal WhatsApp accounts, it blurred the boundary between their home and work, similar to previous findings (Mols and Pridmore, 2021; Wang et al., 2020). Our experts also noted that WhatsApp had a low signal-to-noise ratio, resulting in instances where they missed CataractBot messages: “My WhatsApp has 5000+ [unread] messages… college groups, family groups… I just see messages in the important groups.” (D2). Addressing this issue, the escalation doctor D4 suggested that experts should be recommended to “pin” CataractBot on WhatsApp, ensuring it remains consistently at the top of their chat list.
@@ 6.3.2. Personalization vs Privacy
On the other hand, fueled by the recent hype around AI and LLMs, eight participants expected “personalized” answers to their queries with little concern for privacy, and were slightly disappointed with generic responses. This aligns with prior findings indicating minimal privacy concerns related to digital data in the global south, specifically in India (Okolo et al., 2021). Interestingly, three participants did not even consider their medical data to be “private”, as they stated, “It’s nothing personal to be hidden. All (of us) have common health issues.” (A14). Moreover, two participants questioned why others would be interested in their medical data, as “it is not useful for others” (P18).
Two participants wanted personalization, along with the sharing of patient data, to be an “opt-in” feature where individuals conduct a cost-benefit analysis and make an informed choice based on their preferences. They believed that a personalized bot would offer more relevant, and fewer “I don’t know”, responses.
@@
“If you don’t want to share your information, it (the bot) should say “If you ask personal questions, I won’t have access, so you may have to visit the hospital.”… We have to compromise.” – A21.
@@
Five experts voted to integrate the chatbot with patient data, including schedules and consultation history, to offer personalized responses. D3 suggested integrating patient medical history to facilitate customized responses from the bot and doctors: “We don’t even know whether they’re diabetic, they’re hypertensive. So, when they ask “What should I eat and come?”… I am not sure what to say.” (D3). However, D1 and D2 also voiced skepticism towards the LLMs capability to understand complex patient records accurately. “There are a lot of variables… the surgeon, the patient, their tests… It’s too much information to process.” (D1). In addition, the knowledge base expert cautioned that highly personalized responses would result in minimal updates to the knowledge base, potentially increasing CataractBot-related workload for experts and negatively impacting chatbot adoption among experts.
@@ 7. Discussion
In this paper, we present CataractBot, an LLM-powered experts-in-the-loop WhatsApp chatbot to address the information needs of patients undergoing cataract surgery. Building on limited prior research in multi-stakeholder settings (Yang et al., 2024; Jo et al., 2023), we provide valuable insights into LLM-mediated interactions among patients, attendants, doctors, and coordinators. Our field study revealed positive evidence of CataractBot’s usefulness and usability across stakeholders. Doctor-verified responses were key to patients’ and attendants’ trust in and engagement with the bot. They appreciated that CataractBot instantly answered questions anytime, saving time by reducing the need for hospital calls or visits. Those with limited (tech-)literacy found it easy to use because of the multilingual and multimodal support. Doctors and coordinators commended CataractBot for acting as a facilitator, creating a layer of privacy between them and patients, and providing them the flexibility to verify responses at their convenience.
@@ 7.1. Design Implications
7.1.1. Generalization and Scaling
We designed and evaluated our experts-in-the-loop chatbot system in the context of cataract surgery in India. Expanding it beyond India may require further research. In the United States, for example, insurance plays a crucial role in healthcare. Patients often require prior authorization for treatments, and they would value a bot providing information in navigating these processes (Sun et al., 2013). This necessitates involving insurance-specific experts in knowledge base creation and response verification. Also, our study raised concerns about expert accountability due to the persistence of written communication in the bot. This issue could be more pronounced in the Global North, where doctors are increasingly wary of lawsuit risks (Sun et al., 2013). Balancing these needs and tensions will be critical for CataractBot’s adoption across geographies. Finally, health communication is most effective when users perceive the source as demographically and attitudinally similar to themselves, not just contextually relevant (Kreuter and McClure, 2004). This highlights the need to adapt CataractBot for different settings by (1) parametrizing context-specific details, such as payment methods and hospital protocols, and (2) tailoring language to ensure cultural relevance through familiar terms and metaphors.
Our open-sourced framework can be adapted beyond cataract surgery, to support more complex, atypical, or long-term treatments–e.g., glaucoma surgery, cancer treatment, or pregnancy–where easy access to verified information would be valuable for patients. The applicability can be extended further, to fields beyond the medical domain, such as law, finance and education. In these scenarios, end-users (e.g., defendants, taxpayers, or students) could receive synchronous responses from an LLM using a custom knowledge base, and subsequent asynchronous verified responses from experts (e.g., lawyers, accountants, or teachers). Expert-mediated bots can even be used to upskill less-trained workforce members. For instance, these bots could serve as valuable learning resources for community health workers, who typically have minimal medical training (Yadav et al., 2019).
@@ 7.1.2. Expert-in-the-loop Approaches
Prior studies have raised concerns about the application of LLMs in healthcare scenarios, due to their inconsistency, potential errors, and bias (Denecke et al., 2024; Au Yeung et al., 2023). Moreover, users who are less (technologically) literate face a heightened risk of harm (Kapania et al., 2022). Our experts-in-the-loop framework not only addresses these concerns but also adheres to OpenAI’s usage policies (OpenAI, 2024), which state that “tailored medical/health advice cannot be provided without review by a qualified professional”. By using our framework, chatbots can operate at the “sweet spot of patient-LLM-clinician collaboration” (Hao et al., 2024), where the capabilities of technology and human judgment are integrated. In our study, this ensured reliability and cultivated trust, as information seekers knew that their operating doctor was verifying responses. However, trust in the initial automated answers reduced when these answers were subsequently marked as incorrect/incomplete by the doctor. To address this in future work, one design approach could be to eliminate the initial automated response, providing only expert-verified answers. We caution that this may lead to long wait times, potentially discouraging use. For high-risk use cases like cancer treatment,verified accuracy may be prioritized over immediacy. Future research could compare the two paradigms: a hybrid system with synchronous automated answers followed by asynchronous verification versus a fully asynchronous system delivering only verified answers. Such a study would provide valuable insights into balancing immediate access with the accuracy required for critical medical information in chatbot design.
Doctors are a scare resource in healthcare ecosystems. Our chatbot system can help scale their impact, by reducing the time spent answering routine questions at length during consultations and calls, thus freeing up time for additional patient care. However, as doctors had to verify responses during deployment, CataractBot also added to their workload. To minimize this, one approach is to deliver pre-verified answers to specific questions that the expert has repeatedly marked as correct. Given the potential for LLM hallucinations (Denecke et al., 2024), cached responses must be carefully matched to similar patient demographics and question phrasing before reuse. Another approach is to use a second LLM for quality control while generating each response, as GPT-4’s evaluations of AI-generated responses align well with human expert assessments (Gumma et al., 2024). This maker-checker verification module would enable tracking of the bot’s response quality, especially as the knowledge base expands. The module could also evaluate the confidence level in the generated response, or the perceived urgency of the question, which could be used to prioritize sending nudges to experts for verification.
7.1.3. Multimodality and Personalization
@@
Our system faced challenges in accurately interpreting questions from voice messages due to limitations in transcription and translation technologies. The same inaccurate transcriptions were shared with doctors leading to confusion. Such issues would not arise in deploying a voice-based input modality to a predominantly English-speaking population or a tier-1 language setting in the Global North, as seen in prior work (Yang et al., 2024). Additionally, non-verbal cues in audio messages–such as signs of anxiety, impatience, or the presence of others–were lost in textual transcriptions. This hindered experts’ ability to deliver appropriate care (also noted in (Li et al., 2023)). To address this, we suggest making the original audio clips available to experts on-demand. Beyond audio, considering the visual capabilities of recent LLM models like GPT-4V, such bots could be enhanced to process image-based inputs, aligning with recommendations in prior studies (Jain et al., 2018). For example, a patient could upload a photo of a specific medicine to ask questions about it, instead of providing a written description. While this could improve the user experience for information seekers, it may increase the workload for experts. Further, prior research (Hao et al., 2024) cautions that although LLMs are proficient in text processing, their ability to handle other forms of media, such as image and video, remains less reliable.
@@
Moving on to personalization, while tailored responses can boost adoption and utility by providing more relevant information, they also raise privacy concerns, potentially reducing usage among certain users. This issue may be more pronounced in the developing world, where there is more awareness around privacy. Additionally, experts may face increased workloads as they reference user-specific information for verification, since many personalized edits may not be suitable for inclusion in the generic knowledge base. Low (tech-)literate users might require assistance in making informed decisions about the use of their data (Okolo et al., 2021). A binary approach of either sharing all records or none may not address individual preferences. Instead, users should have the agency to selectively share records they deem relevant and acceptable for the bot’s access. This aligns with the philosophy of India’s National Digital Health Mission (National Health Authority and Family Welfare, 2020). One solution could be to allow users to upload relevant documents or manually enter specific details they want the bot to access. This can empower users to control shared information while benefiting from a personalized experience. Further, when developing CataractBot in May 2023, we chose to use GPT-4, as it was the leading LLM (OpenAI et al., 2023). However, recent open-source models like LLaMA have shown comparable performance to GPT in medical applications (Li et al., 2025). Incorporating these models into the CataractBot system would enable local processing of patient data, helping address privacy concerns.
@@ 7.2. Limitations
We acknowledge several limitations of this work. First, results should be interpreted with caution, given the specificities of the study, including a relatively small sample size, users’ first encounter with a medical chatbot, and the cataract treatment scenario. Positive user responses, minimal increase in expert workload, and other findings should be validated in future studies, preferably with longitudinal designs. Second, system-level analyses, such as scalability testing and power measurement, are needed before deploying the proposed solution more broadly. Finally, although participants reported high trust, repeated “I don’t know” responses or inaccurate information during regular use may erode this trust. We note the potential risk of disseminating inaccurate medical information at scale with chatbot and LLM technologies. Therefore, stronger methods are needed to develop, review, and update the knowledge base, as well as to routinely evaluate the accuracy of bot’s responses.